OpenStudy (anonymous):

Solve by extracting roots: (x+2)^2-7=0

OpenStudy (anonymous):

add 7 to both sides... what equation you have now?

OpenStudy (anonymous):

\[x+2=\pm \sqrt{7}\]

OpenStudy (anonymous):

That's correct. Just need to subtract both sides by 2 to solve for x.

OpenStudy (anonymous):

\[x=\pm \sqrt{7}-2\]

OpenStudy (anonymous):

\[x=-2\pm \sqrt{7}\times \sqrt{-2}\]

OpenStudy (anonymous):

I mean sqrt(-2)...

OpenStudy (anonymous):

I was looking at examples in the book and it says that you are to multiply sqrt (-2) over sqrt (-2)

OpenStudy (anonymous):

That could be for a different problem not really similar to your current problem. Because right now, it is already correct without the multiplication of sqrt(-2).

OpenStudy (anonymous):

Do you know also that \[\sqrt{-2}=2i\] which is am imaginary number.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay then. Glad to be of some help tonight.